# Think Before You Talk: Enhancing Meaningful Dialogue Generation in Full-Duplex Speech Language Models with Planning-Inspired Text Guidance

### v1

@@ 3.2 Text-Guided Full Duplex Dialogue Modeling
After deriving the aligned dialogue turns, the next step is to model double-channel audio data with text guidance. We use GLM-4-Voice (Zeng et al. 2024), an off-the-shelf SLM, as the backbone for training on full-duplex dialogue datasets. GLM-4-Voice includes a speech tokenizer to convert audio into discrete tokens, a language model to process these tokens, and a vocoder to reconstruct speech. Its strong speech-text alignment makes it ideal for the backbone of text-guided speech generation. As illustrated in Figure 3, our text-guided full duplex dialogue modeling framework centers around two interleaving strategies:
@@
Channel-wise Interleaving. Full-duplex dialogue modeling necessitates the simultaneous processing of two-channel audio to capture cross-channel interactions. A common technique to address this is by interleaving the two audio channels into a single sequence, either with token-level interleaving (Wang et al. 2025) or chunk-level interleaving (Veluri et al. 2024; Zhang et al. 2024). We choose chunk-level over token-level interleaving because it demonstrates better performance in our experiments. We first consider channel-wise interleaving for speech tokens. To implement this technique, let $s_{u}$ and $s_{a}$ represent speech tokens from the user and assistant channels. Each channel’s speech tokens are divided into speech chunks with a chunk size of $\tau_{SC}=5$. Given GLM-4-Voice’s frame rate of $fr=12.5\text{Hz}$, each speech chunk corresponds to a 400-ms audio segment. The chunk-interleaving process alternates segments between the two channels, forming a sequence $SC=[SC_{u}^{1},SC_{a}^{1},SC_{u}^{2},SC_{a}^{2},\dots]$. Each speech chunk contains $\tau_{SC}$ speech tokens, with a unique speaker ID token inserted at the beginning to distinguish between the channels. Specifically, $SC_{u}^{i}=[UID,s_{u}^{i\times\tau_{SC}+1},s_{u}^{i\times\tau_{SC}+2},\dots,s_{u}^{i\times\tau_{SC}+\tau_{SC}}]$ and $SC_{a}^{i}=[AID,s_{a}^{i\times\tau_{SC}+1},s_{a}^{i\times\tau_{SC}+2},\dots,s_{a}^{i\times\tau_{SC}+\tau_{SC}}]$ represent the $i$-th user and assistant channel speech chunks, where $UID$ and $AID$ denote the user and assistant ID tokens.
@@
Text-speech Interleaving. We apply the text-guided speech generation strategy on the assistant channel. Specifically, we interleave the text tokens with speech tokens for each turn, starting from the first speech chunk of the turn. Given the text in a turn $W_{j}^{final}$, we first append an “end of text” ($\langle EOT\rangle$) token at the end of it (i.e., $W_{j}^{final}=[t_{j}^{1},t_{j}^{2},...,\langle EOT\rangle]$) and then split it into text chunks $TC_{j}=[TC_{j}^{1},TC_{j}^{2},$...] with chunk size $\tau_{TC}=5$, where $TC_{j}^{i}=[t_{j}^{i\times\tau_{TC}+1},t_{j}^{i\times\tau_{TC}+2},...,t_{j}^{i\times\tau_{TC}+\tau_{TC}}]$ represents the $i$-th text chunk for turn $j$. Second, we calculate the first speech chunk for text-speech interleaving for $W_{j}^{final}$ as
$k_{j}^{start}=\lfloor\frac{\tau_{j}^{start}\times fr}{\tau_{SC}}\rfloor.$ (6)
Then, the interleaving is performed from $k_{j}^{start}$ until the last text chunk as $TC_{j}^{0},SC_{a}^{k_{j}^{start}},SC_{u}^{1},SC_{a}^{k_{j}^{start}+1},$...].
@@
dGSLM (Nguyen et al. 2023)
SyncLLM (Veluri et al. 2024)
Moshi (Défossez et al. 2024)
OmniFlatten (Zhang et al. 2024)
NTPP (Wang et al. 2025)
Wu et al. (Wu et al. 2025)
@@
SLM Semantics Turn-taking (Pearson Correlation)
GPT-score Occ CumDur AvgDur Overall
dGSLM 1.32 0.31 0.24 0.13 0.23
STI 2.55 0.32 0.29 0.14 0.25
SCI 2.71 0.31 0.35 0.17 0.28
Table 2: Comparison of both the semantic (GPT-score, measured on Fisher dataset) and turn-taking abilities (measured on Candor dataset) between our trained baselines and dGSLM. Occ, CumDur, and AvgDur represent occurrence, cumulative duration, and averaged duration, respectively.
@@
SCI 1.81 2.17 2.48 2.71 2.29 1.70 1.87 1.98 2.16 1.93 1.35 1.52 1.67 2.00 1.63 1.95
Moshi TS 1.72 2.06 2.41 2.74 2.23 1.46 1.63 1.82 2.13 1.76 1.23 1.38 1.57 1.96 1.54 1.84
TurnGuide 2.68 2.89 2.93 2.84 2.84 2.08 2.15 2.25 2.26 2.19 2.02 2.12 2.17 2.24 2.14 2.39
TurnGuide L2:1 2.87 3.06 3.08 2.89 2.98 2.20 2.32 2.38 2.35 2.31 2.18 2.33 2.41 2.35 2.32 2.54
TurnGuide L3:1 2.94 3.05 3.08 2.86 2.98 2.25 2.31 2.33 2.32 2.30 2.27 2.38 2.38 2.37 2.35 2.54
Performance Upper Bounds
TurnGuide GT Text 2.61 2.77 2.91 2.82 2.78 2.22 2.32 2.36 2.30 2.30 2.10 2.29 2.30 2.30 2.25 2.44
TextOnly 3.30 3.30 3.28 2.82 3.18 - - - - - - - - - - -
@@
Table 3: Semantic evaluation results (0-5) of different FD-SLMs under different temperature settings by GPT-4o.
@@
4 Experiments
In this section, we assess the effectiveness of TurnGuide, primarily focusing on the semantic meaningfulness of the generated dialogues.
4.1 Experimental Setups
Datasets
We mainly utilize the Fisher dataset (Cieri, Miller, and Walker 2004), a double-channel telephone conversation corpus containing 2,000 hours of audio data, for training and evaluating our FD-SLMs. Each recording in the Fisher dataset spans approximately 10 minutes, which we segment into 120-second clips for training and evaluation. We divide the Fisher dataset into training, validation, and test sets, with detailed information provided in Appendix E. Additionally, we incorporate the Candor dataset (Reece et al. 2023), another double-channel conversation corpus, for evaluation purposes.
Baselines and Our Approach
We consider the following models or methods as our baselines:
  • dGSLM (Nguyen et al. 2023): a dual-Transformer architecture, each modeling an individual audio channel.
  • Speech Token Interleaving (STI): We train the GLM-4-Voice backbone on the Fisher dataset by interleaving the speech tokens from the two channels, one token at a time.
  • Speech Chunk Interleaving (SCI): Similar to Speech Token Interleaving, but it interleaves speech tokens chunk by chunk with a chunk size of 5.
  • Text-speech Token Alignment (Moshi Training Strategy, Moshi TS (Défossez et al. 2024)): The only existing method for text-guided speech generation in e2e FD-SLMs. It involves inserting each text token exactly at the time its corresponding sound is produced. However, Moshi is trained using a wide range of datasets, with Fisher being only one of them, making it not directly comparable to our Fisher-finetuned baselines. To address this, we train the GLM-4-Voice backbone on the Fisher dataset with Moshi’s training methodology to compare it with other baselines. Our detailed implementation of Moshi TS is illustrated in Appendix C.
  • Our Approach (TurnGuide): We train the GLM-4-Voice backbone on the Fisher dataset by utilizing the text-guided full duplex dialogue modeling framework introduced in Section 3.2. The text-speech dataset is pre-processed using the framework illustrated in Section 3.1.
Remark 1
While there are other e2e FD-SLMs apart from dGSLM, most have not been (fully) open-sourced yet, as summarized in Table 1. Therefore, to evaluate our proposed text guidance approach, we train baseline models and show that they perform reasonably well. We then compare TurnGuide against our trained baselines.
Remark 2
Most e2e FD-SLMs demonstrate effectiveness by showing superiority over dGSLM (Veluri et al. 2024; Wang et al. 2025; Défossez et al. 2024). Therefore, we use the same way to demonstrate that our trained baseline performs reasonably well.
Model Training and Inference
The models are trained on the Fisher dataset for 2 epochs on 8 A800 GPUs with a global batch size of 256, a learning rate of 4e-6, a cosine learning rate scheduler, no weight decay, and 20 warmup steps. During inference, we sample 1,000 audio clips from the Fisher and Candor test sets. The first 30 seconds serve as prompts, while the model generates 90-second continuations using two methods: unconditional and conditional generation. In unconditional generation, we train the model with text guidance in both audio channels, and the model generates both channels for the 90-second continuation during inference, whereas in conditional generation, we train the model with only assistant-channel text guidance, and the model generates the assistant-channel audio based on the user-channel audio as context.

### v2

@@ 3.2 Text-Guided Full Duplex Dialogue Modeling Framework
After deriving the aligned dialogue turns, the next step is to model double-channel audio data with text guidance. We use GLM-4-Voice Zeng et al. (2024), an off-the-shelf SLM, as the backbone for training on full-duplex dialogue datasets. GLM-4-Voice includes a speech tokenizer to convert audio into discrete tokens, a language model to process these tokens, and a vocoder to reconstruct speech. Its strong speech-text alignment makes it ideal for the backbone of text-guided speech interaction. As illustrated in the second part of Figure 2, our text-guided full duplex dialogue modeling framework centers around two interleaving strategies:
@@
Channel-wise Interleaving. Full-duplex dialogue modeling necessitates the simultaneous processing of two-channel audio to capture cross-channel interactions. A common technique to address this is by interleaving the two audio channels into a single sequence, either with token-level interleaving Wang et al. (2025) or chunk-level interleaving Veluri et al. (2024); Zhang et al. (2024). We choose chunk-level over token-level interleaving because it demonstrates better performance in our experiments. We first consider channel-wise interleaving for speech tokens. To implement this technique, let $s_{u}$ and $s_{a}$ represent speech tokens from the user and assistant channels. Each channel’s speech tokens are divided into speech chunks with a chunk size of $\tau_{SC}=5$. Given GLM-4-Voice’s frame rate of $fr=12.5\text{Hz}$, each speech chunk corresponds to a 400-ms audio segment. The chunk-interleaving process alternates segments between the two channels, forming a sequence $SC=[SC_{u}^{1},SC_{a}^{1},SC_{u}^{2},SC_{a}^{2},\dots]$. Each speech chunk contains $\tau_{SC}$ speech tokens, with a unique speaker ID token inserted at the beginning to distinguish between the channels. Specifically, $SC_{u}^{i}=[UID,s_{u}^{i\times\tau_{SC}+1},s_{u}^{i\times\tau_{SC}+2},\dots,s_{u}^{i\times\tau_{SC}+\tau_{SC}}]$ and $SC_{a}^{i}=[AID,s_{a}^{i\times\tau_{SC}+1},s_{a}^{i\times\tau_{SC}+2},\dots,s_{a}^{i\times\tau_{SC}+\tau_{SC}}]$ represent the $i$-th user and assistant channel speech chunks, where $UID$ and $AID$ denote the user and assistant ID tokens.
@@
Text-speech Interleaving. We apply text guidance on the assistant channel. Specifically, we interleave the text tokens with speech tokens for each turn, starting from the first speech chunk of the turn.
Given the text in a turn $W_{j}$, we first split it into text chunks $TC_{j}=[TC_{j}^{1},TC_{j}^{2},$...] with chunk size $\tau_{TC}=5$, where $TC_{j}^{k}=[t_{j}^{k\times\tau_{TC}+1},t_{j}^{k\times\tau_{TC}+2},...,t_{j}^{k\times\tau_{TC}+\tau_{TC}}]$ represents the $i$-th text chunk for turn $j$. Then, we append an “end of turn” ($\langle EOT\rangle$) token to the last text chunk, and append an “end of chunk” ($\langle EOC\rangle$) token to the rest of the text chunks. This is to force the model to learn the turn boundary information. Second, we calculate the corresponding speech chunk for text-speech interleaving for each text chunk $TC_{j}^{k}$ in $W_{j}$ as
$k_{j}=\lfloor\frac{\tau_{j}^{k0}\times fr}{\tau_{SC}}\rfloor,$ (6)
where $\tau_{j}^{k0}$ represents the timestamp for the first text token of the $k$-th text chunk in turn $j$. The text-speech interleaving is performed at the $k_{j}$-th chunk of every turn $j$.
@@
dGSLM Nguyen et al. (2023)
SyncLLM Veluri et al. (2024)
Moshi Défossez et al. (2024)
OmniFlatten Zhang et al. (2024)
NTPP Wang et al. (2025)
Wu et al. Wu et al. (2025)
@@
dGSLM 4.70 5.00 4.94 4.03 4.67 - - - - - - - - - - -
STI 4.06 4.94 6.00 7.79 5.70 3.58 4.36 4.94 6.18 4.76 2.76 3.36 3.76 5.24 3.79 4.76
SCI 5.48 6.58 7.52 8.21 6.94 5.15 5.67 6.00 6.55 5.85 4.09 4.61 5.06 6.06 4.94 5.91
Moshi - - - - - 6.70 6.39 5.94 4.21 5.82 6.39 6.42 6.30 4.24 5.85 -
Moshi TS 5.21 6.24 7.30 8.30 6.76 4.42 4.94 5.52 6.45 5.33 3.73 4.18 4.76 5.94 4.67 5.58
TurnGuide 8.12 8.76 8.88 8.61 8.61 6.24 6.58 7.09 7.15 6.76 6.12 6.42 6.58 6.79 6.48 7.27
TurnGuide L2:1 8.70 9.27 9.33 8.76 9.03 6.52 7.03 7.30 7.21 7.03 6.61 7.06 7.30 7.12 7.03 7.70
TurnGuide L3:1 8.91 9.24 9.33 8.67 9.03 6.97 7.24 7.30 7.24 7.18 6.88 7.21 7.21 7.18 7.12 7.79
TurnGuide GT Text 7.91 8.39 8.82 8.55 8.42 6.73 7.03 7.15 6.97 6.97 6.36 6.94 6.97 6.97 6.82 7.39
@@
Table 2: Semantic evaluation results (0-10↑) of different FD-SLMs under different temperature settings by GPT-4o.
4 Experiments
In this section, we assess the effectiveness of TurnGuide, focusing on the semantic meaningfulness and the turn-taking behaviors of the generated dialogues.
4.1 Experimental Setups
4.1.1 Training Data
We utilize the Fisher dataset Cieri et al. (2004), a double-channel telephone conversation corpus containing 2,000 hours of audio data, to train TurnGuide. Following prior works (Nguyen et al., 2023; Veluri et al., 2024), we pre-process each recording in the Fisher dataset into segments of 120-second clips for training and evaluation. We divide the Fisher dataset into training, validation, and test sets, with detailed information provided in Appendix E.
4.1.2 Baselines and Our Approach
We consider the following models or methods as our baselines. 1) dGSLM Nguyen et al. (2023): a dual-Transformer architecture, each modeling an individual audio channel. 2) Speech Token Interleaving (STI): We train the GLM-4-Voice backbone on the Fisher dataset by interleaving the speech tokens from the two channels, one token at a time. 3) Speech Chunk Interleaving (SCI): Similar to Speech Token Interleaving, but it interleaves speech tokens chunk by chunk with a chunk size of 5. 4) Text-speech Token Alignment (Moshi Training Strategy, Moshi TS Défossez et al. (2024)): The only existing method for text-guided speech generation in e2e FD-SLMs. It involves inserting each text token exactly at the time its corresponding sound is produced. However, Moshi is trained using a wide range of datasets, with Fisher being only one of them, making it not directly comparable to our Fisher-finetuned baselines. To address this, we train the GLM-4-Voice backbone on the Fisher dataset with Moshi’s training methodology to compare it with other baselines. Our detailed implementation of Moshi TS is illustrated in Appendix B. 5) Moshi: We also include the original Moshi model for completeness. 6) Our Approach (TurnGuide): We train the GLM-4-Voice backbone on the Fisher dataset by utilizing the text-guided full duplex dialogue modeling framework introduced in Section 3.2. The text-speech dataset is pre-processed using the framework illustrated in Section 3.1.
Remark 1
While there are other e2e FD-SLMs, most have not been (fully) open-sourced yet, as summarized in Table 1. Therefore, we mainly evaluate our approach against dGSLM, Moshi, and our trained baselines.
4.1.3 Model Training and Inference
The models are trained on the Fisher dataset for 2 epochs with a global batch size of 256, a learning rate of 4e-6, a cosine learning rate scheduler, no weight decay, and 20 warmup steps. During inference, we sample 1,000 audio clips from the Fisher test sets. The first 30 seconds serve as prompts, while the model generates 90-second continuations using two methods: unconditional and conditional generation. In unconditional generation, we train the model with text guidance in both audio channels, and the model generates both channels for the 90-second continuation during inference. This setting has potential in two-speaker podcast generation. In conditional generation, we train the model with only assistant-channel text guidance, and the model generates the assistant-channel audio based on the user-channel audio as context, which can be used for spoken interaction purposes.
4.2 Results
We aim to answer three critical Research Questions (RQs) through comprehensive experiments. RQ1: Does our method significantly improve the semantic meaningfulness and coherence of the generated full-duplex dialogue? RQ2: Can our method preserve the natural conversational flow after integrating additional texts into the sequence? RQ3: How does the model perform if we insert text data with different timings and lengths?
Table 3: Full-Duplex-Bench evaluation results. The best performance is bold, and the second-best performance is underlined. Latency is presented in seconds. The scores in brackets indicate temperature.
Dimension Pause Handling Backchannel Smooth Turn Taking User Interruption
Data Candor ICC Candor Synthetic
Metric TOR ($\downarrow$) TOR ($\downarrow$) Freq ($\uparrow$) JSD ($\downarrow$) TOR ($\uparrow$) Latency ($\downarrow$) TOR ($\uparrow$) GPT-4o($\uparrow$) Latency ($\downarrow$)
dGSLM 0.935 0.691 0.015 0.934 0.975 0.352 0.917 0.201 2.531
Moshi 0.980 1.000 0.001 0.957 0.941 0.265 1.000 0.765 0.257
TurnGuide L2:1 (0.8) 0.390 0.667 0.013 0.927 0.773 0.567 0.558 1.349 2.263
TurnGuide L2:1 (0.9) 0.441 0.519 0.019 0.860 0.773 0.357 0.689 1.441 1.870
TurnGuide L2:1 (1.0) 0.587 0.630 0.021 0.887 0.807 0.282 0.866 1.471 1.566
TurnGuide L2:1 (1.3) 0.789 0.500 0.043 0.814 0.950 0.102 1.000 1.150 0.900
@@ 4.2 Results
4.2.1 RQ1
To evaluate the semantics of the generated dialogue, the most intuitive method is to conduct human evaluations. However, considering the time, cost, and potential inconsistencies associated with human evaluations, we propose utilizing the GPT-score. Specifically, given a generated double-channel dialogue audio, we first transcribe the audio from each channel into text with sentence-level timestamps. For this process, we employ the stable-ts package Jian (2023) with the Whisper-large-v3 model. Subsequently, the transcribed sentences from both channels are grouped based on their start times, effectively representing the dialogue between the two speakers in text form along the timeline. Finally, the consolidated text is analyzed using GPT-4o Hurst et al. (2024), which assesses the meaningfulness and coherence of the dialogue on a scale of 0 to 5 with detailed rubrics. The complete GPT-4o prompts for unconditional and conditional generation settings are shown in Appendix G. To demonstrate the reliability of GPT-based evaluation, we conduct human evaluations, showing strong consistency between human- and GPT-evaluated scores (see Appendix A).
We demonstrate the effectiveness of TurnGuide through evaluations in both unconditional and conditional settings. In the “Conditional - All” scenario, GPT-4o uses ordered transcripts from both channels to assess responses based on user speech. However, we observe that GPT-4o overly prioritizes user-assistant interactions, neglecting the coherence of assistant-generated speech within its channel. To address this, the “Conditional - Assistant” setting provides only the assistant-channel transcript, enabling a focused analysis of semantic quality within the assistant channel. We also observe that GPT-4o tends to assign conservative scores for this task, which is evidenced by the ground truth text from the Fisher dataset achieving an average GPT-score of 3.3. Therefore, we rescale all GPT-score results from 0-3.3 to 0-10.0.
We compare TurnGuide with various baselines under different temperatures (as shown in Table 2) and make the following observations. 1) Significant performance improvements over baselines. TurnGuide performs significantly better than SCI, Moshi TS (over 30% performance gain) and Moshi (over 20% performance gain), showing that TurnGuide successfully leverages the text information to boost the meaningfulness and coherence of the generated dialogue. 2) Speech-only methods and text-speech methods excel at different temperatures. SCI consistently performs the best at the largest temperature (1.3), whereas TurnGuide tend to perform the best at temperature 1.0. Additionally, TurnGuide shows more performance gain under lower temperatures, such as 0.8 and 0.9. 3) Focusing more on text generation makes the model perform better. In “TurnGuide L2:1" and “TurnGuide L2:1" settings, we set the training loss weight of text and speech tokens to be 2:1 and 3:1, and the results show that they consistently outperform the standard 1:1 setting (“TurnGuide"). 4) TurnGuide achieves performance comparable to its upper bounds. In Table 2, “TurnGuide GT Text" refers to the GLM-4-Voice model trained with our proposed framework but switches to the ground truth transcriptions from the Fisher dataset. TurnGuide performs similarly with TurnGuide GT Text in the 1:1 loss setting (7.27 vs. 7.39), and the TurnGuide 2:1 and 3:1 outperform the upper bound in 1:1 loss.
Further validation with perplexity. To facilitate the semantic evaluation using GPT-score, perplexity is employed to further demonstrate the superior semantic abilities of TurnGuide. Specifically, we calculate perplexity using the same ordered transcript used in the GPT-score evaluation. Perplexity is measured with three distinct language models: Fisher-text-only, GLM-4-Voice, and Llama-3.1-8b-Instruct Dubey et al. (2024). The Fisher-text-only model, trained using Fisher-provided text transcript data on the GLM-4-Voice base model, serves as the ground truth for text-based dialogue distribution. GLM-4-Voice acts as the SLM backbone, while Llama-3.1-8b-Instruct serves as the text-only model backbone. We analyze the perplexity of generated dialogues across different temperature settings in both conditional and unconditional scenarios. As illustrated in Figure 3, TurnGuide consistently achieves lower perplexity compared to baseline approaches, demonstrating superior semantic modeling abilities.
4.2.2 RQ2
Evaluating the naturalness of FD-SLM interaction involves assessing various turn-taking events in the full-duplex dialogues they generate. Traditional FD-SLM studies (Nguyen et al., 2023; Veluri et al., 2024; Défossez et al., 2024) often analyze turn-taking behaviors at the corpus level, comparing the model-generated statistics to the corresponding ground-truth statistics. However, Lin et al. (2025) highlights that corpus-level statistics are often difficult to interpret, making them less suitable as an evaluation method. To overcome this limitation, they introduce the Full-Duplex-Bench benchmark, designed to evaluate the fine-grained turn-taking behaviors of FD-SLMs. This benchmark specifically assesses core turn-taking abilities, such as Smooth Turn-taking, User Interruption, Pause Handling, and Backchanneling. The detailed explanation of the four turn-taking events is in Appendix H.
We compare TurnGuide trained with 2:1 loss with the baseline approaches on Full-Duplex-Bench (as shown in 3) and highlight the following. 1) TurnGuide demonstrates superior interaction naturalness compared to baseline approaches. TurnGuide surpasses baselines on nearly all the metrics of the four turn-taking events, except for the TOR of smooth turn-taking and the latency of user interruption. 2) TurnGuide achieves optimal turn-taking performance when the temperature is set to 1.3. At this temperature setting, TurnGuide outperforms not only the baselines but also TurnGuide at other temperatures.
SLM Conditional - All
0.8 0.9 1.0 1.3 AVG
TurnGuide 6.24 6.58 7.09 7.15 6.76
Too Long 5.73 5.97 6.30 6.52 6.12
Too Short 6.06 6.52 6.73 6.70 6.52
Too Early 4.88 5.06 5.64 6.88 5.61
Too Late 5.27 5.76 6.06 6.70 5.94
Table 4: Semantic evaluation results of TurnGuide on different text insertion timings and lengths.
4.2.3 RQ3
We aim to support the claim that the text insertion in the full duplex dialogue modeling process exhibits the insertion timing and length challenges. Specifically, we simulate the following four text insertion scenarios and demonstrate the superiority of our approach. 1) Too Long: Each text chunk contains 20 text tokens. 2) Too Short: Each text chunk contains 2 text tokens. 3) Too Early: The insertion time of every text chunk is shifted to 4s earlier. 4) Too Late: The insertion time of every text chunk is shifted to 4s later.
The results presented in Table 4 highlight the following. 1) Text insertion timings and lengths indeed pose challenges to the full duplex dialogue modeling process. We observe that all the simulated settings significantly degrade the model performance. 2) Model performance is particularly sensitive to insertion timing. The Too Early and Too Late scenarios result in greater performance degradation compared to Too Long and Too Short.
5 Conclusion
We present TurnGuide, a novel method to improve the dialogue generation abilities of e2e FD-SLMs. By segmenting assistant speech into turns and training the model with channel-wise interleaving and turn-level text-speech interleaving strategies, our approach effectively addresses key challenges of text insertion timing and length, ensuring coherent and natural dialogue generation. Experiments show significant improvements of our method in semantic quality and various natural turn-taking events, validating its effectiveness for spoken dialogue systems and paving the way for more meaningful and human-like speech interactions.
Limitations
While the paper introduces TurnGuide, a novel dynamic turn-level text-speech interleaved generation approach for enhancing e2e FD-SLMs, it still presents several limitations. First, the effectiveness of TurnGuide is primarily validated on the Fisher dataset, which, although widely used, may not fully capture the diversity and complexity of real-world conversational scenarios since Fisher is a telephone conversation dataset. However, addressing this limitation would necessitate the effort of open-sourcing real-world double-channel dialogue datasets that encompass a broader range of styles. Second, the evaluation relies heavily on GPT-4o-based automated metrics (GPT-score) for semantic assessment, which, despite being aligned with human judgments in most cases, may not perfectly reflect the nuanced human perceptions of dialogue quality.
Appendix A Judgement Alignment between Human and GPT-4o in Semantic Evaluation
Difference 1 2 3 Overall
Accuracy 7/10 9/10 9/10 25/30
SLMs M-MOS N-MOS
SCI 2.24 2.47
Moshi 2.69 2.75
TurnGuide 3.26 3.23
Ground Truth 4.25 4.29
Table 6: Absolute rating results of the human evaluations assessing the Meaningfulness (M-MOS) and Naturalness (N-MOS) of the dialogues.
In this paper, we primarily utilize GPT-4o to evaluate the semantics of the generated dialogues. This raises a crucial question: Can AI evaluation scores accurately capture human preferences? To investigate this, we conduct two human listening tests, including pairwise comparison and absolute rating, to examine the alignment between GPT-evaluated scores and human judgments.
A.1 Pairwise Comparison
Pairwise comparison involves organizing the generated dialogues into pairs and having human judges assess which dialogue in each pair exhibits better semantic quality. This experiment aims to determine whether GPT-4o can effectively differentiate between good and bad dialogues. Specifically, we randomly select 30 pairs of generated dialogues that vary in their GPT-scores: 10 pairs with a GPT-score difference of 1, 10 pairs with a difference of 2, and 10 pairs with a difference of 3. To ensure unbiased evaluation, the order of dialogues within each pair is randomized. Additionally, judges are provided with both the audio of the dialogues and their transcripts, generated using Whisper-large-v3 with the stable-ts package. We recruit three human judges, all of whom have native English proficiency, to carry out the evaluations. All human raters provided informed consent before participating and were informed that their annotations may be used for academic research and reported in an academic publication. The complete instructions given to the participants are presented in Figure 9.
A.2 Absolute Rating
Absolute rating involves human evaluators directly assessing the quality of dialogues generated by various models using a 5-point Likert scale. In this method, we can gather the human-evaluated scores and compare them with the GPT-scores in Table 2. A strong alignment between humans and AI is observed when the ranking of scores for each model remains consistent across both human and AI evaluations. We randomly choose 10 audio prompts and obtain their continuations by SCI, Moshi, TurnGuide, and the Ground Truth, under each model’s best-performing temperature. This results in a total of 40 audio samples. Following the exact settings and instructions outlined in Veluri et al. (2024), human evaluators rate the Meaningfulness (M-MOS) and Naturalness (N-MOS) of the dialogues. Same as pairwise comparison, judges are provided with both the audio of the dialogues and their transcripts. We recruit nine human judges with native English proficiency for this task.
The results presented in Table 6 also demonstrate a strong alignment between human and AI evaluations. Notably, the ranking of the models aligns with the AI’s assessments in Table 2, further validating the reliability of our automated semantic evaluation approach. Moreover, among all the models, TurnGuide achieves the highest scores in both meaningfulness and naturalness, underscoring its effectiveness as a superior general-purpose FD-SLM.
Appendix B Moshi TS Implementation
First, Moshi utilizes eight codebooks, where every time step includes eight speech tokens for each channel. We adapt this technique for the GLM-4-Voice backbone, which relies on a single codebook method, allotting one speech token per channel per time step. Then, the tokens from each channel are interleaved at each time step, making Moshi rely on a token-level interleaving approach. In the original implementation of Moshi, a text-channel token precedes each assistant-channel speech token, which could either be a genuine text token, a text pad token, or an end-of-text pad token. Moshi also employs text-speech interleaving at token-level, where each text token is incorporated at the precise moment it is spoken. Consequently, if a time step features an audible text token, the corresponding text token is inserted. If not, either a text pad token or an end-of-text pad token is inserted, determined by whether the subsequent text-channel token is an actual text token. We follow the original Moshi implementation by extracting the timestamp of each word with Whisper-timestamped package Louradour (2023) and the medium Whisper model Radford et al. (2023). The illustration of our implemented Moshi TS is shown in Figure 4.